Question 1. 2cosx^4 -3cosx^2 +1=0
Question 2. 4sinx^4 +3sinx^2 -1=0
Question 3. Cos3x= (sqrt(3)/2)
Question 4. 2cosx^2 -2sinx^2=1
Question 5. 2sinxcosx=1
Please, solve all the five equations above, and I require the answer in between 0 degrees and 360 degrees [0,360].

Answer by Alan3354(69443)

Question 1. 2cosx^4 -3cosx^2 +1=0
-----------
(2cos^2 - 1)*(cos^2 - 1) = 0
cos^2 = 1/2, etc
cos^2 = 1, etc
============================
#2 is similar
#3 is easy
#4 is similar
----------------
Question 5. 2sinxcosx=1
sin(2x) = 1
2x = 90 degs